# Robust image representations with counterfactual contrastive learning

### v1

@@ Results
Figure 3: Pneumonia detection results with linear probing (frozen encoder, solid lines) and finetuning (unfrozen encoder, dashed lines) for models trained with the SimCLR objective. Results are reported as average ROC-AUC over 3 seeds, shaded areas denote +/- one standard error. We also compare to a supervised baseline initialised with ImageNet weights. CF-SimCLR consistently outperforms encoders trained with standard SimCLR and SimCLR+ where counterfactuals are added to the training set for linear probing and performs best overall for full model finetuning.
@@
Figure 4: Breast density results with linear probing (frozen encoder, solid lines) and finetuning (unfrozen encoder, dashed lines) for models trained with SimCLR. Results are reported as average one-versus-rest macro ROC-AUC over 3 seeds, shaded areas denote +/- one standard error. CF-SimCLR overall performs best across ID and OOD data, improvements are largest in the low data regime and on under-represented scanners.
@@
Figure 7: t-SNE projections of embeddings from 16,000 randomly sampled test images from mammography encoders trained with SimCLR, SimCLR+ and CF-SimCLR. Encoders trained with SimCLR and SimCLR+ exhibit domain clustering. CF-SimCLR embeddings are substantially less domain-separated and the only disjoint cluster exclusively contains breasts with implants, semantically different. Thumbnails show a randomly sampled image from each ‘implant’ cluster. Adapted from [11].
In this section, we compare the representations quality and robustness for various pre-training paradigms. First, standard SimCLR. Secondly, SimCLR+ where we train a model using classic SimCLR on a training set enriched with domain counterfactuals. Finally, CF-SimCLR combining SimCLR with our proposed counterfactual contrastive pair generation framework. We then repeat the same analysis for models pre-trained with the DINO objective, comparing DINO, CF-DINO and DINO+. Note that in SimCLR+ (resp. DINO+) counterfactuals and real images are not paired during the contrastive learning step, they are all considered as independent training samples. As such SimCLR+/DINO+ represent the common paradigm of simply enriching the training set with synthetic examples. In CF-SimCLR/CF-DINO, on the other hand, we systematically pair real images with their corresponding counterfactual for positive pair creation (Fig. 1). We compare the effect of these three pretraining strategies on chest X-rays and mammograms. For chest X-rays, we evaluate the quality of the learned representations by assessing downstream performance on pneumonia detection. For mammography, we focus on the task of breast density prediction (important for risk modelling).
Does counterfactual contrastive learning improve performance and robustness under acquisition shift?
Results on pneumonia detection (Fig. 3) show that CF-SimCLR outperforms the SimCLR baseline (orange versus blue), across datasets, irrespective of the amount of labels, with improvements particularly striking for linear probing. The largest performance gains can be observed in low levels of labels, on the ID scanner under-represented during training (Imaging) and the external datasets. Mammography results in Fig. 4 also show that CF-SimCLR consistently outperforms the SimCLR baseline across most ID scanners for both linear probing and finetuning, particularly when the amount of labelled data is limited (<20k). On the external VinDR dataset, CF-SimCLR beats both baselines on both scanners in the low data regime. Crucially, CF-SimCLR pretraining particularly benefits scanners under-represented in the training set (all except Selenia Dimensions for EMBED and PlanMed Nuance for VinDr), regardless of the amount of labelled data. For both modalities, the improvement on external datasets is particularly worth highlighting given that the encoder was not exposed to these external domains during pretraining (nor during counterfactual generation).
Crucially, performance improvements and increased robustness to acquisition shift equally hold for encoders trained with the DINO-v2 objective, demonstrating the versatility of the proposed method. In Fig. 6, we can see that CF-DINO outperforms DINO for all EMBED scanners, across all levels of labels, for both linear probing and full model finetuning. Again the gains mostly affect scanners under-represented during training with CF-DINO closing the performance gap between the majority scanner (Selenia Dimensions) and the other scanners. As for CF-SimCLR, similar observations are made for the VinDr minority scanner with CF-DINO substantially improving performance on PlanMed nuance for low levels of labels. For chest X-rays, in Fig. 6, we can see that training with CF-DINO closes the performance gap between both PadChest scanners for both linear probing evaluation and finetuning. Indeed, without CF-DINO there is a substantial performance drop from Phillips images to Imaging images, whereas this gap is much smaller with performance improving substantially with CF-DINO across all levels of labels for images from Imaging. Similarly, CF-DINO outperforms DINO on the external RSNA Pneumonia dataset by a substantial margin. For CheXpert, CF-DINO outperforms DINO in model finetuning (and linear probing with 10% of labels), however, it slightly underperforms on the 25% and 100% for linear probing. Note that results on the CheXpert dataset are to be interpreted with some caution as labels are NLP sourced and in general of rather low quality [33] which may explain the consequent drop of performance for all models between CheXpert and the expert-labelled RSNA Pneumonia dataset.
What is the benefit of counterfactual contrastive learning over simply extending the training set with the generated counterfactual data?
We have shown that the models trained with counterfactual contrastive learning outperform models trained with standard contrastive pairs. However, it is important to note that models trained with CF-SimCLR (resp. CF-DINO), are exposed to additionally (synthetic) data during training. Hence, in this section, to isolate the effect of the ‘smart pair creation’ proposed in this work, we compare it to another baseline, SimCLR+ (resp. DINO+), where we use the same extended training set but where generated samples are considered as independent samples and are not paired with real images during training. Overall, CF-SimCLR outperformed SimCLR+ consistently across all experimental settings. Similarly, CF-DINO outperformed DINO+ across all EMBED datasets and both performed similarly on VinDR. On chest X-rays, CF-DINO outperformed DINO+ by a large margin on PadChest and RSNA Pneumonia for linear probing, and both performed similarly for model finetuning. We especially noticed that the performance gains of SimCLR+ (resp. DINO+) over SimCLR were not very stable, improving for some domains while performing on par with standard SimCLR (resp. DINO) for others (compare orange versus green in Figs. 3 and 4 for SimCLR and Figs. 6 and 6 for DINO). In general, the counterfactual contrastive approaches offered more consistent performance improvements. The starkest differences were observed on scanners under-represented during training (see Fig. 4). These experimental results are in-line with theoretical benefits of counterfactual contrastive learning: by explicitly creating realistic cross-domains positive pairs during training, we directly encourage the network to create domain-agnostic image representations. The more domain-agnostic image representations are, the bigger the expected improvement in terms of robustness to acquisition shift. Such an increase in robustness leads to higher performance on devices under-represented during training, even more so in limited data settings. The increase in domain alignment is directly visible in the t-SNE plots of feature embeddings in Fig. 7, where embeddings of models trained with SimCLR and SimCLR+ exhibit clear domain clustering, whereas the model trained with CF-SimCLR generate embeddings where domains are less separated.
Computational considerations
@@
To implement counterfactual contrastive learning, we need to train an additional causal image generation model, this raises the question of computational overhead. Fortunately, the HVAE used here to generate image counterfactuals is relatively lightweight (as opposed to alternative generative approaches such as diffusion models). This HVAE not only trains relatively fast (only 20 epochs needed), it is also relatively frugal in terms of VRAM requirements (20 GB of GPU VRAM were sufficient), which has the advantage of low hardware requirements. In terms of generation speed, we were able to generate over 1 million 224 x 224 mammography images in under 7 hours (on an NVIDIA RTX-3090 GPU). These computational requirements need to be put in perspective compared to the requirements of the contrastive pretraining step. Contrastive pretraining is resource intensive both in terms of VRAM (large batch sizes) and in terms of training time. Each SimCLR model was trained for 500 epochs and required 2x46GB VRAM. For DINO, memory requirements were even higher (6x46GB VRAM for a batch size of 300). The overhead of the counterfactual generation part is negligible, and its benefits clearly outweigh the added computational costs.
Robustness beyond acquisition shifts
The main focus of this work is to improve robustness to acquisition shift. However, by simply changing the types of generated counterfactuals, the counterfactual contrastive pair generation framework can be easily extended to other distribution shifts. For example, we may use the same framework to increase robustness to population shift, improving performance of under-represented subgroups. We illustrate this on chest radiography in Fig. 8, where we generate sex counterfactuals instead of domain counterfactuals in the positive pair creation step, aiming to improve group-wise performance of the ‘male’ and ‘female’ subgroups. We can see that the baseline approach (SimCLR) performs sub-optimally on female patients for pneumonia detection across all datasets. Using cross-subgroups counterfactuals positive pairs by intervening on biological sex, we observe performance improvements across subgroups in all datasets and for most levels of training labels.
@@
Figure 8: Improving sub-group performance with counterfactual contrastive learning. Pneumonia detection results with linear probing for encoders trained with SimCLR and SexCF-SimCLR. In SexCF-SimCLR we generate sex counterfactuals instead of domain counterfactuals for positive pair generation to improve robustness to subgroup shift and ultimately performance on under-represented subgroups. Results are reported as average ROC-AUC over 3 seeds, shaded areas denote +/- standard error. Sex CF-SimCLR performs best overall ID and OOD, increasing performance for both gender subgroups.
Discussion and Conclusions
@@
In this work, we present counterfactual contrastive learning, a novel contrastive pair generation framework enhancing the robustness of contrastively-learned image representations to domain shifts. Evaluating across five datasets, two modalities and two clinically-relevant classification tasks, we show that the proposed counterfactual contrastive pretraining approach yields higher downstream performance than standard contrastive pretraining; improvements which are particularly noticeable for domains less represented in the training set.Moreover, the proposed counterfactual pair generation method is agnostic to the choice of the contrastive objective, as demonstrated by our experiments showing that counterfactual positive pair generation improves results for models using SimCLR as well as DINO-v2 objectives. Importantly, we show that CF-SimCLR (resp. CF-DINO) also improve performance on external domains, not included in the pretraining set compared to SimCLR (resp. DINO). Our experiment on subgroup counterfactual contrastive learning demonstrates its wider applicability beyond acquisition shifts.
@@
Naturally, gains arising from counterfactual contrastive learning are bounded by the ability to generate realistic domain changes. Generated counterfactuals must be of sufficient quality to faithfully capture the variation relevant for pre-training. Our results demonstrate that current counterfactual image generation models can already produce images of sufficient quality to significantly improve learned representation over the baseline approaches. Note that our proposed approach is compatible with other generative-based augmentation methods such as generating additional images for under-represented classes or subgroups [37, 38], as we can also apply the counterfactual generation to synthetic images. Our experiments comparing SimCLR+ with CF-SimCLR show that the proposed approach has benefits beyond complementing the training set with synthetic data and that the proposed counterfactual pair generation framework fundamentally changes the organisation of the embedding space (Fig. 7), as such it may complement other approaches to enhance training set diversity.
Acknowledgments

### v2

@@ 5.1 Does counterfactual contrastive learning improve performance and robustness under acquisition shift?
Figure 5: Pneumonia detection results with linear probing (frozen encoder, solid lines) and finetuning (unfrozen encoder, dashed lines) for models trained with the SimCLR objective. Results are reported as average ROC-AUC over 3 seeds, shaded areas denote +/- one standard error. We also compare self-supervised encoders to a supervised baseline initialised with ImageNet weights.
Figure 6: ROC-AUC difference to SimCLR baseline for CF-SimCLR and SimCLR+ for pneumonia detection. The top row depicts results with linear probing, bottom row shows results with model finetuning. Results are reported as average ROC-AUC difference compared to the baseline (SimCLR) over 3 seeds, error bars denote +/- one standard error. CF-SimCLR consistently outperforms encoders trained with standard SimCLR and SimCLR+ (where counterfactuals are added to the training set) for linear probing, and performs best overall for full model finetuning.
@@
Figure 7: Breast density results with linear probing (frozen encoder, solid lines) and finetuning (unfrozen encoder, dashed lines) for models trained with SimCLR. Results are reported as average one-versus-rest macro ROC-AUC over 3 seeds, shaded areas denote +/- one standard error. CF-SimCLR performs best overall across ID and OOD data, and improvements are largest in the low data regime and on under-represented scanners.
Figure 8: ROC-AUC difference between SimCLR and CF-SimCLR (resp. SimCLR+) for breast density assessment. The top two rows denote results with linear probing, and the bottom two rows show results with model finetuning. Results are reported as average macro ROC-AUC difference compared to the baseline (SimCLR) over 3 seeds, error bars denote +/- one standard error. CF-SimCLR overall performs best across ID and OOD data, improvements are largest in the low data regime and on under-represented scanners.
@@
Figure 10: ROC-AUC difference between DINO and CF-DINO (resp. DINO+). Top two rows denote results with linear probing, bottom two rows results with model finetuning. Results are reported as average macro ROC-AUC difference compared to the baseline (DINO) over 3 seeds, error bars denote +/- one standard error. CF-DINO overall performs best across ID and OOD data, improvements are largest in the low data regime and on under-represented scanners.
@@
Figure 12: ROC-AUC difference to DINO baseline for CF-DINO and DINO+ for pneumonia detection. The top row depicts results with linear probing, bottom row show results with model finetuning. Results are reported as average ROC-AUC difference compared to the baseline (DINO) over 3 seeds, error bars denote +/- one standard error.
@@
Figure 13: t-SNE projections of embeddings from 16,000 randomly sampled test images from mammography encoders trained with SimCLR, SimCLR+ and CF-SimCLR. Encoders trained with SimCLR and SimCLR+ exhibit domain clustering. CF-SimCLR embeddings are substantially less domain-separated and the only disjoint cluster exclusively contains breasts with implants, semantically different. Thumbnails show a randomly sampled image from each ‘implant’ cluster. Adapted from (Roschewitz et al., 2024).
5.3 Computational considerations
@@
To implement counterfactual contrastive learning, we need to train an additional causal image generation model, this raises the question of computational overhead. Fortunately, the HVAE used here to generate image counterfactuals is relatively lightweight (as opposed to alternative generative approaches such as diffusion models). This HVAE not only trains relatively fast (e.g. only 20 epochs needed for EMBED, about 250k steps), it is also relatively frugal in terms of VRAM requirements (20GB of GPU VRAM were sufficient), which has the advantage of low hardware requirements. In terms of generation speed, we were able to generate over 1 million 224x224 mammography images in under 7 hours (on an NVIDIA RTX-3090 GPU). These computational requirements need to be put in perspective compared to the requirements of the contrastive pretraining step. Contrastive pretraining is resource intensive both in terms of VRAM (large batch sizes) and in terms of training time. Each SimCLR model was trained for 450 epochs for EMBED (resp. 1,000 epochs for PadChest), and required 2x46GB VRAM. For DINO, memory requirements were even higher (6x46GB VRAM for a batch size of 300). The overhead of the counterfactual generation part is negligible, and its benefits clearly outweigh the added computational costs.
5.4 Ablation study on the impact of counterfactual quality on downstream performance
In this section, we aim to assess the effect of the counterfactual image generation model quality on the downstream performance of the proposed counterfactual contrastive objective. To this end, we here compare three counterfactual image generation models, trained on EMBED:
  • HVAE-: an HVAE-based counterfactual image generation as per Ribeiro et al. (2023) trained for one epoch only (13k steps).
  • HVAE: the HVAE used in the experiment so far, trained for 20 epochs, without counterfactual finetuning.
  • HVAE+FT: where we additionally add the counterfactual finetuning step, as proposed by Ribeiro et al. (2023), to the HVAE model. This step allows for improved effectiveness by further finetuning the counterfactual generation models using guidance from external classifiers.
First, we assess counterfactual quality for all three models using the metrics proposed in Monteiro et al. (2022). Results in Fig. 14 show that effectiveness varies strongly across models, especially for scanners under-represented in the training set (e.g. Senographe Pristina). Moreover, results in Table 2 show that while effectiveness varies strongly across the compared models, reversibility and composition are strong for all models, indicating that all models preserve identity well. This can also be seen in the qualitative examples in Fig. 15. Strong semantic identity preservation is key for creating meaningful contrastive pairs, as the network will consider any changes across views in a positive pair as non-semantic information to disregard. Besides computational considerations, this further motivates the use of the HVAE counterfactual approach in this study, over, for example, diffusion-based counterfactual generation approaches as diffusion models have been shown to not always preserve identity well (Preechakul et al., 2022; Mokady et al., 2023).
Figure 14: Effectiveness comparison for the three counterfactual models considered in this ablation study, by intervention. Computed on 8,304 validation set samples.
CF model $\mathrm{Effectiveness}$ $\mathrm{Reversibility}^{(1)}$ $\mathrm{Composition}^{(1)}$
do(scanner)
HVAE- 49% 0.001 8e-12
HVAE 77% 0.002 5e-12
HVAE+FT 96% 0.004 5e-12
Table 2: Axiomatic soundness metrics (Monteiro et al., 2022) for the three models considered in this ablation study. Computed on the 8,304 validation set samples.
Figure 15: Qualitative of comparison of the three counterfactual generation models, HVAE-, HVAE and HVAE+FT compared in the ablation study. For each model we show generated counterfactuals as well as direct effect maps. Direct effects give a visual depiction of the increase in effectiveness across the three models from top to bottom. We also observe that all models preserve semantic identity very well, a key aspect in positive pair creation contrastive learning.
Figure 16: Effect of counterfactual quality on downstream performance. Results are reported as average macro ROC-AUC difference compared to the baseline (SimCLR) over 3 seeds for linear probing, error bars denote +/- one standard error. We compare running CF-SimCLR with (i) HVAE- a counterfactual generation model of lesser effectiveness, (ii) HVAE the generation model used in the rest of this study, (iii) HVAE+FT a counterfactual generation model with higher effectiveness.
In Fig. 16, we compare the downstream performance of the three CF-SimCLR encoders trained using generated counterfactuals from the three counterfactual generation models analysed above: HVAE-, HVAE and HVAE+FT. The graphs depict differences in ROC performance compared to the SimCLR baseline (i.e. no counterfactuals) across several levels on labels and scanners. These results first show that even when using a model with relatively low effectiveness, CF-SimCLR with HVAE- is either on par with or outperforms the baseline (positive performance differences), especially on under-represented scanners and for lower levels of labels. However, as effectiveness increases, the downstream performance for under-represented scanners tends to increase. We can see that the effectiveness of the HVAE model is substantially better compared to the HVAE- model for generating Lorad Selenia (+30%) and Senographe Pristina (+70%) counterfactuals. Further improving effectiveness with counterfactual finetuning (i.e. HVAE+FT) leads to striking improvements in the most under-represented scanner Senograph Pristina (where effectiveness increases by 60% between HVAE and HVAE+FT). On this scanner, we see improvements of 4%, 2%, 0.5% and 1.8% ROC-AUC between CF-SimCLR with HVAE and CF-SimCLR with HVAE+FT across the various levels of labels, while also observing noticeable improvements on ID Lorad Selenia and the OOD VinDr Mammomat scanner. However, we also do notice somewhat diminishing returns of effectiveness improvements, where performance on ClearviewCSM, for example, is not substantially changing between CF-SimCLR with HVAE-, HVAEand HVAE+FT despite effectiveness improving from 85% to 97%, with similar results observed on Senograph 2000D where increase in effectiveness did not help to improve downstream performance. Results indicate that a scanner effectiveness above 80% is satisfactory for CF-SimCLR.
5.5 Robustness beyond acquisition shifts
@@
Figure 17: Improving sub-group performance with counterfactual contrastive learning. Pneumonia detection results with linear probing for encoders trained with SimCLR and SexCF-SimCLR. In SexCF-SimCLR, we generate sex counterfactuals instead of domain counterfactuals for positive pair generation to improve robustness to subgroup shift and, ultimately, performance on under-represented subgroups. Top row: performance for the male (solid line) and female (dashed line) subgroups, reported as average ROC-AUC over 3 seeds, shaded areas denote +/- standard error. Bottom row: performance disparities across the two subgroups, reported as average ROC-AUC difference between the male and female subgroups, over 3 seeds. Sex CF-SimCLR reduces sub-group disparities for all datasets, substantially increasing performance on the female sub-group when limited amounts of labels are available, both on ID and OOD datasets.
The main focus of this work is to improve robustness to acquisition shift. However, by simply changing the types of generated counterfactuals, the counterfactual contrastive pair generation framework can be easily extended to other distribution shifts. For example, we may use the same framework to increase robustness to population shift, improving the performance of under-performing subgroups. We illustrate this on chest radiography in Fig. 17, where we generate sex counterfactuals instead of domain counterfactuals in the positive pair creation step, aiming to reduce performance disparities between the ‘male’ and ‘female’ subgroups. We can see that the baseline approach (SimCLR) performs sub-optimally on female patients for pneumonia detection across all datasets (top row in Fig. 17). Using cross-subgroup counterfactuals positive pairs by intervening on biological sex, we observe performance improvements across subgroups in all datasets and for most levels of training labels. For example, on PadChest we observe an improvement of 2% ROC-AUC with 3249 labelled samples for females compared to the SimCLR baseline (resp. 1% with 16247 labels). On CheXpert, we also get +5% ROC-AUC with 1381 labelled samples, and +3% with 3452 labels. For PadChest and CheXpert, we even observe small performance improvements on the male subgroup. On RSNA however, while Sex-CFSimCLR successfully reduces the performance gap across both groups (bottom row), this here comes at the cost of a slight performance reduction in the male subgroup (top row), a phenomenon commonly known as the ‘levelling down’ effect in the fairness literature (Zietlow et al., 2022). Overall, analysing performance differences across subgroups (bottom row in Fig. 17), we notice that Sex-SimCLR yields a consistent reduction in performance disparities across all datasets. Like in previous experiments, we notice that performance improvements are most notable in settings where it is most beneficial to have aligned representations across subgroups: in limited label settings and for the underperforming subgroups.
6 Discussion and Conclusions
@@
In this work, we present counterfactual contrastive learning, a novel contrastive pair generation framework enhancing the robustness of contrastively-learned image representations to domain shifts. Evaluating across five datasets, two modalities and two clinically-relevant classification tasks, we show that the proposed counterfactual contrastive pretraining approach yields higher downstream performance than standard contrastive pretraining, improvements which are particularly noticeable for domains less represented in the training set. Moreover, the proposed counterfactual pair generation method is agnostic to the choice of the contrastive objective, as demonstrated by our experiments showing that counterfactual positive pair generation improves results for models using SimCLR as well as DINO-v2 objectives. Importantly, we show that CF-SimCLR (resp. CF-DINO) also improves performance on external domains, which are not included in the pretraining set compared to SimCLR (resp. DINO). Our experiment on subgroup counterfactual contrastive learning demonstrates its broader applicability beyond acquisition shifts.
In general, improvements arising from counterfactual contrastive learning are most visible in settings where learning domain-aligned representations presents a strong advantage over having domain-clustered representations (as learned in standard SimCLR). This is notably the case for scanners under-represented during finetuning, and in settings with limited label availability, where insufficient data is present to learn a ‘scanner-wise’ decision boundary. Conversely, when a large amount of labels is available for model finetuning (e.g. up to 223k labelled samples on EMBED and 65k on PadChest), all pre-training strategies often converge to similar performance. This is explained by the fact that, with high amounts of labels, the benefit of domain-aligned representations is reduced as enough data is present at finetuning time to learn meaningful decision boundaries for each scanner, even if representations are clustered per domain. Moreover, it is important to highlight that, in the case of full model finetuning, representations may change substantially over the pre-trained representations, further explaining why downstream performance differences are often more visible in linear probing experiments. However, it is important to highlight that the counterfactual contrastive objective performs consistently either better or, depending on the number of labels, on par with existing baselines.
@@
Naturally, gains arising from counterfactual contrastive learning are bounded by the ability to generate realistic domain changes. Generated counterfactuals must be of sufficient quality to capture the variation relevant for pre-training faithfully. Our results demonstrate that current counterfactual image generation models can already produce images of sufficient quality to significantly improve learned representation over the baseline approaches. Importantly, results in our ablation study in Section 5.4 show that even when the effectiveness of the counterfactual image generation model is moderate, CF-SimCLR remains on par with, or still outperforms the SimCLR baseline. However, as effectiveness increases, the performance on particularly under-represented scanners can substantially increase (e.g. CF-SimCLR with HVAE+FT on the Pristina scanner). As image synthesis models continue to improve, we may expect further improvements with counterfactual contrastive learning in future works.
Importantly, the choice of the intervention variable in the counterfactual image generation step depends on which types of shifts downstream models should be robust against. In the main experiments (Section 5.1), where the goal was to achieve scanner robustness, we intervened on the scanner variable. In the experiment in Section 5.5, the goal was to be robust against gender shift. Hence, we generated sex counterfactuals. A combination of multiple variables could also be considered in future work. In practice, we can summarise the process to determine which variables to intervene on, as follows:
  1. 1.
    First, an in-depth data analysis should be performed to determine likely causes of biases and subgroup disparities, and to determine which variable the downstream representations should be robust to, i.e. should ignore.
  2. 2.
    The set of identified variables in step 1, should then be discussed with domain experts to decide which variables downstream models should be robust against, and whether any interactions between these variables should be considered in the causal graph used to generate the counterfactuals.
  3. 3.
    Train the counterfactual generation model with the previously identified causal graph.
  4. 4.
    Use the generated counterfactuals to create cross-domain pairs in the counterfactual contrastive framework, with interventions randomly sampled among all possible values when sampling positive pairs.
Note that our proposed approach is compatible with other generative-based augmentation methods, such as generating additional images for under-represented classes or subgroups (Ktena et al., 2023; Khosravi et al., 2024), as we can also apply the counterfactual generation to synthetic images. Our experiments comparing SimCLR+ with CF-SimCLR show that the proposed approach has benefits beyond complementing the training set with synthetic data and that the proposed counterfactual pair generation framework fundamentally changes the organisation of the embedding space (Fig. 13), as such it may complement other approaches to enhance training set diversity.
Data and code declaration
All datasets used in this study are publicly available. PadChest (Bustos et al., 2020) can be found at https://bimcv.cipf.es/bimcv-projects/padchest/, the RSNA Pneumonia Detection Dataset (Stein et al., 2018) can be downloaded at https://www.kaggle.com/c/rsna-pneumonia-detection-challenge, CheXpert (Irvin et al., 2019) can be found at https://stanfordaimi.azurewebsites.net/datasets/8cbd9ed4-2eb9-4565-affc-111cf4f7ebe2. We only used the publicly available part of the EMBED dataset (Jeong et al., 2023) instructions can be found at https://github.com/Emory-HITI/EMBED_Open_Data/tree/main, the VinDR Mammo (Nguyen et al., 2023) can be found at https://vindr.ai/datasets/mammo. All the code required to reproduce our experiments is publicly available at https://github.com/biomedia-mira/counterfactual-contrastive.
Author contributions
M.R. and B.G. conceived and designed the experiments. M.R. and F.D.S. wrote the code. M.R. performed the experiments. All authors analysed results and wrote the paper.
Acknowledgements